# Latency-Aware Generative Semantic Communications with Pre-Trained Diffusion Models

### v1

@@ I Introduction
In this paper, we develop a universal generative semantic communications framework with pre-trained foundation models utilizing stable diffusion. We claim two key benefits in comparison with the existing semantic communication frameworks. Firstly, the foundation models possess a vast general knowledge leveraging the intensive self-supervised training process on huge amount of data. This alleviates the need for a shared knowledge base/graph between the semantic transmitter and receiver, obviating the need for corresponding knowledge sharing overheads imposed in current SemCom frameworks [13, 14, 15]. This vast general knowledge also makes the proposed Generative SemCom framework applicable to various datasets and tasks thereby achieving universality. Secondly, the adoption of pre-trained models allows a separation-based SemCom architecture, alleviating the need for end-to-end joint training of the transmitter and receiver, which is required in many SemCom frameworks [13, 14, 15]. Such a separation-based architecture offers better compatibility with the existing design of wireless communication networks in comparison with the end-to-end methods. The contributions of this work are three-fold:
@@
The rest of this paper is organized as follows. In Section II we provide our proposed generative semantic communication framework with foundation models and present the corresponding semantic and latency-aware adaptive semantic communication scheme. Simulation results are provided in Section III and we conclude the paper in Section IV.
Notations: Boldface lower and upper-case symbols denote column vectors and matrices, respectively. $[K]$ denotes the set $\{1,2,...,K\}$. For a vector ${\bf v}$, $\left|{\bf v}\right|$ denotes the length of the vector, ${\bf v}\cdot{\bf v}$ denotes the dot production.
@@ II-A Multi-modal Semantic Decomposition and Synthesis
The decomposed multi-modal semantics, i.e., prompt and conditioning signals, are then compressed, coded, and modulated, and transmitted over the wireless channel. At the receiver, after demodulation, decoding, and decompression, the received multi-modal semantics are simultaneously fed into a pre-trained generative model for high-fidelity synthesis of the source signal. We note that state-of-the-art generative models, e.g. stable diffusion (SD) [4], allow generation using only the prompt, which results in a relatively high distortion. On the other hand, some semantic information content of the source signal can be more efficiently communicated using other modalities. The conditioning signals improve the semantic fidelity as well as distortion at additional communication costs.
@@ II-B1 Re-transmission-based Communication of the Textual Prompt
The data size for the prompt is typically several times smaller than the other semantic modalities, we can thus perfectly afford several re-transmissions of the prompt while the other semantic modalities are being transmitted over orthogonal channels. On the other hand, the prompt contains the most important semantic contents of the source signal. Based on this, we propose a cyclic redundancy check (CRC)-based re-transmission mechanisms for the prompt to guarantee its reliable communication. Assuming the prompt is transmitted in packets of length $L$ bits, the packet error rate (PER) for prompt is given by
@@
where $k$ and $b$ are parameters determined by the channel coding scheme and the modulation order [16]. We denote the coding rate and the modulation order as $r$ and $M_{0}$, respectively. Hence, the expected total transmission delay for the prompt is given by
@@ II-B2 Adaptive Modulation and Coding for other Semantic Modalities
For the transmission of the other semantic modalities, we adopt an adaptive MQAM modulation and coding scheme to cope with varying wireless channel. With such a scheme, the achievable rate can be generally expressed as [17]
@@
where $\alpha$ and $\beta$ are parameters determined by the coding schemes, $\text{BER}_{i}$ denotes the bit error rate (BER) of the $i$-th conditioning signal. Thereby, the transmission delay of $i$-th semantic stream is ${T_{i}}(p_{i},\text{BER}_{i})=\frac{{\left|{{{\bf{v}}_{i}}}\right|}}{{{R_{i}}(p_{i},\text{BER}_{i})}}$, $\forall i\in[I]$.
@@ III-A System setup
We consider a generative image semantic communication setting, where the receiver is interested in the main objects inside the image, e.g. “car”, “building”, and its general structure, i.e. object shapes and locations. In this context, we consider a two-modal semantic decomposition framework, including a textual prompt to convey the main objects, and an edge map as the conditioning signal to convey the general image structure. To extract the prompt, we adopt the GPT-4 model [22] in this work [^2] Alternative smaller image-to-text models, e.g. BLIP [23], Oscar [24], UNIMO [25], etc. are also possible.. We also adopt the holistically-nested edge detection model [26], to extract a gray-scale edge map from the original image at the transmitter. The prompt is encoded to bits using UFT-8 [27], and the edge map is compressed to bits using learned nonlinear transform coding [28]. At the receiver, we adopt the state-of-the-art pre-trained SD model [4] for conditional image generation from the received prompt and the edge map. The wireless transmission parameters for the two semantic streams are listed in Table I.
@@
Figure 2: Normalized CLIP and MS-SSIM versus BER of the edge map.
@@ III-B Semantic quality metrics
Various reference-based or no-reference metrics can be adopted to evaluate the quality of the synthesized signal based on the semantic communication intent. In this work, we adopt two metrics, one to measure the semantic similarity, and the other to measure the structural similarity between the original and the synthesized images at the Gen SemCom receiver. In order to assess the quality of the most important semantic information conveyed by the prompt, we use the contrastive language-image pretraining (CLIP) model [21]. Specifically, we define the CLIP metric as the normalized cosine similarity between the CLIP embedding of the original and synthesized images, i.e. $f({\bf x})$ and $f(\hat{\bf x})$, given by $\text{CLIP}=(\frac{f({\bf x})\cdot f(\hat{\bf x})}{\|f({\bf x})\|_{2}\cdot\|f(\hat{\bf x})\|_{2}}+1)/2$. For structural similarity, we use the multi-scale structural similarity (MS-SSIM) [18] metric.
@@ III-C Monotonicity of the quality metrics
The end to end performance of the proposed framework depends on behaviour of the pre-trained stable diffusion model at the receiver, i.e. its sensitivity to the BER of the conditioning signal, and the requirements of the semantic quality metrics. Hence, we investigate monotonicity of the CLIP and MS-SSIM metrics to apply Lemma 1 for latency-aware adaptive semantic communications. Specifically, through intensive simulations in presence of random bit errors in the proposed framework, we demonstrate that the normalized CLIP/MS-SSIM metrics are monotonically non-increasing functions of the BER of the edge map, as depicted in Fig. 2. The normalized CLIP and MS-SSIM are defined as $\frac{\text{CLIP}}{\text{CLIP}_{0}}$ and $\frac{\text{MS-SSIM}}{\text{MS-SSIM}_{0}}$, respectively, where CLIP0 and MS-SSIM0 are the reference values of these metrics for perfect transmission of the edge map. We note that “CLIP, Prompt + Edge Map” surpasses “CLIP, Prompt Only” when $\text{BER}\leq 10^{-4}$, indicating that incorporating the edge map enhances the semantic quality as interpreted by the SD model. On the contrary, at higher BER levels of the edge map, specifically when $\text{BER}>10^{-3}$, the edge map inaccuracies become detrimental, causing the SD model to misinterpret the textual semantics. Moreover, transmitting the edge map alongside the prompt invariably boosts the MS-SSIM in comparison to relying solely on the prompt.
Figure 3: Visualization of the semantic quality of our proposed framework on a natural image. The bpp of prompt and edge map is 0.0024 and 0.017.
@@ III-D Visual quality of the proposed framework
In Fig. 3, we provide illustrations to demonstrate the visual quality of our proposed framework on a sample natural image. These results demonstrate good semantic quality of the proposed framework even at moderate BER values. These results are for transmission rates significantly lower than the typical values in current networks, i.e. bit per pixel (bpp) $\ll 0.1$.
@@ III-E Latency-aware Adaptive Semantic Communication
We evaluate the optimal semantic latency in terms of the average SNR, i.e. $\overline{\gamma}=\frac{p_{\text{T}}\mathbb{E}(|h_{0}|^{2})}{(B_{0}+B_{1})N_{0}}$, as a channel quality indicator. With the investigation in subsection III-C, we apply Lemma 1 and solve the resulting nonlinear system of equations numerically. Fig. 4 quantifies the semantic quality-latency trade-off for the proposed Gen SemCom framework, and illustrates the optimal wireless parameters versus average SNR $\overline{\gamma}$ at various target semantic quality (CLIP, MS-SSIM) and BER values. It is evident that to achieve a higher target semantic quality (i.e. higher CLIP/MS-SSIM and lower BER) one should allocate more power to the transmission of the edge map, which in turn, increases the latency. Furthermore, when the channel SNR deteriorates, the sender should use a lower modulation order for the edge map to maintain the semantic quality at a target acceptable level. This will also increase the expected number of re-transmissions for the prompt.
For performance comparison, we adopt a semantic-unaware transmission benchmark, in which the prompt and edge map are treated equally and transmitted as one data stream regardless of the unequal importance of each semantic modality. For a fair comparison, the maximum power $p_{\rm T}$ and bandwidth $B_{0}+B_{1}$ are allocated for the transmission of this single stream with adaptive MQAM employed. The prompt BER becomes the bottleneck for single-stream transmission, and hence, the BER is kept at $10^{-7}$ to ensure reliable reception of the prompt. From Fig. 4(b), the proposed semantic-aware multi-stream transmission significantly reduces the transmission latency at the lower SNR region, i.e., $\overline{\gamma}<12$ dB.
Finally, in Table II, we provide modulation adaptation guidelines to achieve various semantic qualities (CLIP, MS-SSIM) in varying channel qualities. Here, $M_{1}$ is the modulation order for the edge map, and $T$ is the transmission latency.
Figure 4: Optimal wireless parameters versus average SNR $\overline{\gamma}$ at various target BERs: (a) Percentage of power for prompt; (b) Transmission Latency; (c) Bits/symbol for edge map; (d) Average numbers of prompt re-transmissions.

### v2

@@ I Introduction
In this paper, we develop a universal generative semantic communications framework with pre-trained foundation models. We claim two key benefits in comparison with the existing semantic communication frameworks. Firstly, the foundation models possess a vast general knowledge leveraging the intensive self-supervised training process on huge amount of data. This alleviates the need for a shared knowledge base/graph between the semantic transmitter and receiver, obviating the need for corresponding knowledge sharing overheads imposed in current SemCom frameworks [14, 16, 17, 15, 18]. This vast general knowledge also makes the proposed Generative SemCom framework applicable to various datasets and tasks thereby achieving universality. Secondly, the adoption of pre-trained models allows a separation-based SemCom architecture, alleviating the need for end-to-end joint training of the transmitter and receiver, which is required in many SemCom frameworks [14, 16, 17, 15]. Such a separation-based architecture offers better compatibility with the existing design of wireless communication networks in comparison with the end-to-end methods. The proposed framework is specifically suitable for scenarios which require communication of huge multi-modal data with stringent latency and reliability requirements, e.g. the wireless metaverse, extended/mixed reality (XR/MR), holographic teleportation, and the internet of senses. The contributions of this work are three-fold:
@@
Notations: Boldface lower and upper-case symbols denote column vectors and matrices, respectively. $[K]$ denotes $\{1,2,...,K\}$. $[{\bf v}]_{n}$ and $\left|{\bf v}\right|$ denote the $n$-th element and the length of vector ${\bf v}$, respectively. Finally, ${\bf v}\cdot{\bf v}$ is the dot production, and $Pr\{\cdot\}$ denotes probability of an event.
@@ II-A Multi-modal Semantic Decomposition and Synthesis
The decomposed multi-modal semantics, i.e., prompt and conditioning signals, are then compressed, coded, and modulated, and transmitted over the wireless channel. At the receiver, after demodulation, decoding, and decompression, the received multi-modal semantics are simultaneously fed into a pre-trained generative model for high-fidelity synthesis of the source signal. We note that state-of-the-art generative models, e.g. stable diffusion (SD) [5], allow generation using only the prompt, which results in a relatively high distortion. On the other hand, some semantic information content of the source signal can be more efficiently communicated using other modalities. The conditioning signals improve the semantic fidelity as well as distortion at additional communication costs. Finally, the desired dimensions of the signal to be synthesized by GenAI are assumed to be fixed and known to the receiver.
@@ II-B1 Re-transmission-based Communication of the Textual Prompt
The textual prompt is a super compact message that is highly sensitive to errors. Even a single bit error can change a character/word causing a significant semantic error. Thereby, our framework uses re-transmissions if the prompt is received with any error. Based on this, we propose a cyclic redundancy check (CRC)-based re-transmission mechanism for the prompt to guarantee its reliable communication. Assuming the prompt is transmitted in packets of length $L$ bits, the packet error rate (PER) can be expressed as [19]
@@
where $k$ and $b$ are parameters determined by the channel coding scheme and the modulation order. We denote the coding rate and the modulation order as $r$ and $M_{0}$, respectively. Hence, the expected total transmission delay for the prompt is given by
@@ II-B2 Adaptive Modulation and Coding for other Semantic Modalities
For the transmission of the other semantic modalities, we adopt an adaptive MQAM modulation and coding scheme to cope with varying wireless channel. Denote the bit error rate (BER) of the $i$-th, $i\in[I]$, conditioning signal as $\text{BER}_{i}=\frac{1}{|{\bf{v}}_{i}|}\sum\nolimits_{n=1}^{|{\bf{v}}_{i}|}Pr\{[\hat{\bf{v}}_{i}]_{n}\neq[{\bf{v}}_{i}]_{n}\}$, where $\hat{\bf{v}}_{i}$ denotes the corresponding received data stream of the conditioning signal. With such a scheme, the achievable rate can be generally expressed as [20]
@@
where $\alpha$ and $\beta$ are parameters determined by the coding schemes. Thereby, the transmission delay of $i$-th semantic stream is ${T_{i}}(p_{i},\text{BER}_{i})=\frac{{\left|{{{\bf{v}}_{i}}}\right|}}{{{R_{i}}(p_{i},\text{BER}_{i})}}$, $\forall i\in[I]$.
@@ III-A System setup
We consider a generative image semantic communication setting, where the receiver is interested in the main objects inside the image, e.g. “car”, “building”, and its general structure, i.e. object shapes and locations. In this context, we consider a two-modal semantic decomposition framework, including a textual prompt to convey the main objects, and an edge map as the conditioning signal to convey the general image structure. To extract the prompt, we adopt the GPT-4 model [25] in this work. We also adopt the holistically-nested edge detection model [29], to extract a gray-scale edge map from the original image at the transmitter. The prompt is encoded to bits using UFT-8 [30], and the edge map is compressed to bits using learned nonlinear transform coding [31]. At the receiver, we adopt the state-of-the-art pre-trained SD model [5] for conditional image generation from the received prompt and the edge map. The wireless transmission parameters for the two semantic streams are listed in Table I.
@@
Figure 2: Normalized CLIP and MS-SSIM versus BER of the edge map: (a) Prompt generated by GPT-4; (b) Prompt generated by BLIP.
@@ III-B Semantic quality metrics
Various reference-based or no-reference metrics can be adopted to evaluate the quality of the synthesized signal based on the semantic communication intent. In this work, we adopt two metrics, one to measure the semantic similarity, and the other to measure the structural similarity between the original and the synthesized images at the Gen SemCom receiver. In order to assess the quality of the most important semantic information conveyed by the prompt, we use the contrastive language-image pretraining (CLIP) model [24]. Specifically, we define the CLIP metric as the cosine similarity between the CLIP embedding of the original and synthesized images, i.e. $f({\bf x})$ and $f(\hat{\bf x})$, given by $\text{CLIP}=(\frac{f({\bf x})\cdot f(\hat{\bf x})}{\|f({\bf x})\|_{2}\cdot\|f(\hat{\bf x})\|_{2}}+1)/2$. For structural similarity, we use the multi-scale structural similarity (MS-SSIM) [21] metric.
@@ III-C Monotonicity of the quality metrics
The end-to-end performance of the proposed framework depends on behaviour of the pre-trained GenAI model at the receiver, i.e. its sensitivity to the BER of the conditioning signal, and the requirements of the semantic quality metrics. Hence, we investigate monotonicity of the CLIP and MS-SSIM metrics to apply Lemma 1 for latency-aware adaptive SemCom. Specifically, through intensive simulations in presence of random bit errors in the proposed framework, we demonstrate that the normalized CLIP/MS-SSIM metrics are monotonically non-increasing functions of the BER of the edge map, as depicted in Fig. 2. Note that the proposed framework works independent of the choice of the AI models adopted at the transmitter and receiver. For example, the GPT-4 model can be replaced by smaller image-to-text models, e.g. BLIP [26], Oscar [27], UNIMO [28], at the cost of a slight degradation of the semantic quality metrics. These smaller models can be implemented locally on-device. The absolute CLIP values achieved with GPT-4 and BLIP, under error-free transmission of the edge map, are $0.918$ and $0.896$, respectively, while the resulting MS-SSIM values are similar. The normalized CLIP and MS-SSIM are defined as $\frac{\text{CLIP}}{\text{CLIP}_{0}}$ and $\frac{\text{MS-SSIM}}{\text{MS-SSIM}_{0}}$, respectively, where CLIP0 and MS-SSIM0 are the reference values of these metrics for perfect transmission of the edge map. We note that “CLIP, Prompt + Edge Map” surpasses “CLIP, Prompt Only” when $\text{BER}\leq 10^{-4}$, indicating that transmitting the edge map enhances the semantic quality as interpreted by the SD model. On the contrary, when $\text{BER}>10^{-3}$, the inaccuracies in the received edge map due to the communication errors become detrimental, causing the SD model to misinterpret the prompt semantics. For example in Fig. 3, the “fence” is missing in the generated image due to the inaccuracies of the received edge map at $\text{BER}=10^{-3}$, despite the word “fence” being part of the prompt. Moreover, transmitting the edge map alongside the prompt invariably boosts the MS-SSIM in comparison to relying solely on prompt.
Figure 3: Visualization of the semantic quality of our proposed framework. The absolute CLIP/MS-SSIM values are reported in this figure for comparison.
@@ III-D Visual quality of the proposed framework
In Fig. 3, we provide illustrations to demonstrate the visual quality of our proposed framework on a sample natural image. These results demonstrate a good semantic quality for the proposed framework for ultra-low-rate transmission at bit per pixel (bpp) values as low as $0.0024$ and $0.017$ for the prompt and the edge map, respectively.
@@ III-E Latency-aware Adaptive Semantic Communication
We evaluate the optimal semantic latency in terms of the average SNR, i.e. $\overline{\gamma}=\frac{p_{\text{T}}\mathbb{E}(|h_{0}|^{2})}{(B_{0}+B_{1})N_{0}}$, as a channel quality indicator. With the investigation in subsection III-C, we apply Lemma 1 and solve the resulting nonlinear system of equations numerically. Fig. 4 quantifies the semantic quality-latency trade-off for the proposed Gen SemCom framework, and illustrates the optimal wireless parameters versus average SNR $\overline{\gamma}$ at various target semantic quality, i.e. normalized (CLIP, MS-SSIM), and BER values. It is evident that to achieve a higher target semantic quality (i.e. higher CLIP/MS-SSIM and lower BER) one should allocate more power to the transmission of the edge map, which in turn, increases the latency. Furthermore, when the channel SNR deteriorates, the sender should use a lower modulation order for the edge map to maintain the semantic quality at a target acceptable level. This will also increase the expected number of re-transmissions for the prompt.
For performance comparison, we adopt a semantic-unaware transmission benchmark, in which the prompt and edge map are treated equally and transmitted as one data stream regardless of the unequal importance of each semantic modality. For a fair comparison, the maximum power $p_{\rm T}$ and bandwidth $B_{0}+B_{1}$ are allocated for the transmission of this single stream with adaptive MQAM employed. The prompt BER becomes the bottleneck for single-stream transmission, and hence, the BER is kept at $10^{-7}$ to ensure reliable reception of the prompt. From Fig. 4(b), the proposed semantic-aware multi-stream transmission significantly reduces the transmission latency at the lower SNR region, i.e., $\overline{\gamma}<12$ dB. Finally, in Table II, we provide modulation adaptation guidelines to achieve various semantic qualities (CLIP, MS-SSIM) in varying channel qualities. Here, $M_{1}$ is the modulation order for the edge map, and $T$ is the transmission latency.
Figure 4: Optimal wireless parameters versus average SNR $\overline{\gamma}$ at various target BERs: (a) Percentage of power for prompt; (b) Transmission Latency; (c) Modulation order (bits per symbol) for the edge map; (d) Average numbers of prompt re-transmissions. The (CLIP, MS-SSIM) notation in the legend represents the normalized CLIP/MS-SSIM values achieved for each curve.
@@
III-F Computation latency
The computation latency depends on the choice of the AI models and the implementation scenario. For the on-device implementation scenario, smaller models like BLIP (55 GFLOPs) [26] and mobile diffusion (153 GFLOPs) [32] are used. Using a state-of-the-art A17 Pro chip (35 TFLOPS/s), the resulting computation latencies are 1.6 ms for BLIP and 4.6 ms for mobile diffusion. These are comparable to the transmission latencies as reported in Fig. 4(b) and TABLE II, emphasizing the need to minimize the transmission latency for reduced end-to-end latency.